OpenStudy (anonymous):

3x+3y=3 (solving for x)

OpenStudy (anonymous):

3x+3y = 3 -3y 3x = 3 - 3y 3x/3 = (3-3y)/3 x = 1- y hope i helped :)

OpenStudy (anonymous):

3x+3y = 3 -3y 3x = 3 - 3y 3x/3 = (3-3y)/3 x = 1- y

OpenStudy (anonymous):

YES! thanks :)

OpenStudy (anonymous):

You're very welcome:)